1. Let T: R2 – R? be the map reflection in the line y = x—you may assume this T is linear, let Eº be the standard basis ofa) Use the equations (1) - + (1) + (-) and (9) - : (I) - (-to find the matrix ola de: e) Use your previous answers to verify
Answer #1

Reflection about line yax T(4) - (9 ) )1 1 - 1 (11) = ( :. ( 15) - (1) (ا, الا ( (211 T ( 2 ) = ( 1 / 2 ) (CI ،1) . (ارم) - 5

☺. I(949) = (029) I (1,1)=(1,0) I (1,-1) = (11-1) (try Etho + 4 Coral (/) = (1,0) +!(0.1) (1,-1) = (10)-1(011) e) - 4(llo eva

@ I (04) = 6c,y) identity transformation 5(1,0) = (1,0) ICO. :(0.1) ; (,0) = a (1,1) + 6(1,5-) (1,0) =(a+b, a-b) a+b=1 a-bao